OpenStudy (savannah_noelle):

Can someone show me how to work these out please? I'll give a medal <3 1. -17y=da solve for d 2. 3a-p=r solve for a 3. -6d+5b=n solve for d 4. J=F(1+rt) solve for t 5. z=9+y, for y;z=5 y= 6.A certain rule for medication expresses a relationship between the recommended dosages for a child and for and for an adult. The rule states that a child's dosage C equals the product of the child's weight W in pounds divided by 140 and the adult dosage A. Write a formula for C in terms of W and A. Then solve this formula for A. The formula for C in terms of W and A is C=

OpenStudy (anonymous):

yo

OpenStudy (savannah_noelle):

@zendaya-replay-guy can you help?

OpenStudy (anonymous):

well....... i can probably do it on a calculator

OpenStudy (anonymous):

hey @Darius25 and @johnweldon1993 can u help here?

OpenStudy (johnweldon1993):

I'll work out a couple and see if that helps you get going.. Well it will just be using algebra to isolate a variable... 1. -17y=da solve for d So to isolate d on one side...you need to divide everything by 'a' -17y d = ----- a 2. 3a-p=r solve for a Same thing here...add 'p' to both sides of the equation 3a = r + p Now divide everything by 3 to isolate 'a' a = r + p ------ 3 Make sense?

OpenStudy (anonymous):

@Savannah_Noelle

OpenStudy (savannah_noelle):

Okay so for number one the equation would be D/A=-17Y/A Does the "A" cancel out?

OpenStudy (anonymous):

u got this? @johnweldon1993

OpenStudy (johnweldon1993):

\[\large \frac{D \cancel{A}}{ \cancel{A}} = \frac{-17y}{A} \] \[\large D = \frac{-17y}{A}\] So yes the A cancels

OpenStudy (savannah_noelle):

Okay thank you, is that the final answer?

OpenStudy (johnweldon1993):

That would be your final answer for 1) yes

OpenStudy (savannah_noelle):

@johnweldon1993 @zendaya-replay-guy Thank you guys so much! :) but I have a question. How do you know when you need to divide, multiply, add, etc for thee equations?

OpenStudy (savannah_noelle):

Also, the website i'm on says that a=r+p/3 was the wrong answer

OpenStudy (johnweldon1993):

Alright...so it's all about inverses of operations...Lets look again at 1) We had \[\large -17y = da\] We want to isolate D....well we can see that D is being MULTIPLIED by A ...so in order to cancel that out...we need to DIVIDE by A....

OpenStudy (johnweldon1993):

Hint* a = (r + p)/3 Use parenthesis...it makes all the difference

OpenStudy (savannah_noelle):

Ohh okay, that makes sense

OpenStudy (johnweldon1993):

if that is how you put your answer...you entered \[\large a = r + \frac{p}{3}\] But the answer is \[\large a = \frac{r + p}{3}\]

OpenStudy (savannah_noelle):

Okay so for number 3: -6d+5b=n solve for d, I should solve by dividing everything by n?

OpenStudy (johnweldon1993):

No...The goal would be to isolate d because that is what you are solving for... -6d + 5b = n You want this to look like -6d = ...something... Make sense? you want 'd' to be alone on 1 side of the equation... So you would start by subtracting 5b from both sides of this equation -6d + 5b = n - 5b -5b ---------------- -6d = n - 5b With me so far?

OpenStudy (savannah_noelle):

Yes

OpenStudy (johnweldon1993):

Alright....so now...we want to solve for 'd' not -6d.... So since 'd' is being MULTIPLIED by -6....to cancel that we...?

OpenStudy (savannah_noelle):

divide by -6

OpenStudy (johnweldon1993):

Right...but remember divide EVERYTHING by -6 -6d = n- 5b ----------- -6 -6 d = n - 5b ------ -6

OpenStudy (savannah_noelle):

Okay thank you :) So for number 4: J=F(1+rt) solve for t, would I use the distributive property and get J=F1+FRT?

OpenStudy (johnweldon1993):

Right...now you want to isolate FRT first...so what would you do first?

OpenStudy (savannah_noelle):

subtract?

OpenStudy (johnweldon1993):

Correct...subtract what?

OpenStudy (savannah_noelle):

F1, so now I have J-F1=FRT

OpenStudy (johnweldon1993):

Just a short thing.. \[\large F1 = F\] it is just F being multiplied by 1 which is just F :) But yes correct!

OpenStudy (savannah_noelle):

Oh okay :) and so now I need to divide?

OpenStudy (johnweldon1993):

Right! divide what?

OpenStudy (savannah_noelle):

everything by FR?

OpenStudy (johnweldon1993):

Perfect :) so t = ?

OpenStudy (savannah_noelle):

T=J-F/FR

OpenStudy (johnweldon1993):

Right...just remember parenthesis t = (J - F)/(FR)

OpenStudy (savannah_noelle):

Thank you so much! :)

OpenStudy (johnweldon1993):

Anytime :) think you got the last 2?

OpenStudy (savannah_noelle):

I thought number 5 would equal -4 since 5 substitutes for z, but it's saying that's wrong :/

OpenStudy (savannah_noelle):

@johnweldon1993

OpenStudy (johnweldon1993):

Hmm...well z = 9 + y First it says solve for 'y' so you would subtract 9 from both sides... y = z - 9 When z = 5 y = ? y = z - 9 y = 5 - 9 y = -4 You are correct...but maybe they just want the equation? y = z - 9?

OpenStudy (savannah_noelle):

Yes I figured that out, and I'm now on number 6. I figured out that the formula for C is C=(w/140)A But now it's asking what is the formula when solved for A? I don't know how to do that

OpenStudy (johnweldon1993):

Well look how you have it... You want to solve for 'A' C = (w/140)A Looks like A is just being MULTIPLIED by something right? so divide by w/140 HINT* when you divide by a fraction....you flip the fraction and turn it into multiplication...ie \[\huge \frac{C}{\frac{w}{140}} = \frac{\frac{w}{140}}{\frac{w}{140}}A\] Will become \[\huge \frac{140}{w}C = A\] Make sense? All I did was flip the fraction...and then turned it into multiplication

OpenStudy (savannah_noelle):

Oh okay I was typing in (W/140)C

OpenStudy (savannah_noelle):

Thank you for everything! I really appreciate it :)

OpenStudy (johnweldon1993):

Anytime :)